OpenStudy (anonymous):

i typed it wrong, the question is 5 1/6 = x -7/8. thank you

OpenStudy (anonymous):

\[5{1\over6} = -{7\over8}x\]\[{31\over6} = {-7x \over 8}\]

OpenStudy (anonymous):

cross multiplication: (-7x)*(6) = (8)*(31) -42x = 248 x = 248/(-42)

OpenStudy (anonymous):

thank you